### v1

# ISLES’24: Improving final infarct prediction in ischemic stroke using multimodal imaging and clinical data

Abstract
Accurate estimation of core (irreversibly damaged tissue) and penumbra (salvageable tissue) volumes is essential for ischemic stroke treatment decisions. Perfusion CT, the clinical standard, estimates these volumes but is affected by variations in deconvolution algorithms, implementations, and thresholds. Core tissue expands over time, with growth rates influenced by thrombus location, collateral circulation, and inherent patient-specific factors. Understanding this tissue growth is crucial for determining the need to transfer patients to comprehensive stroke centers, predicting the benefits of additional reperfusion attempts during mechanical thrombectomy, and forecasting final clinical outcomes. This work presents the ISLES’24 challenge, which addresses final post-treatment stroke infarct prediction from pre-interventional acute stroke imaging and clinical data. ISLES’24 establishes a unique 360-degree setting where all feasibly accessible clinical data are available for participants, including full CT acute stroke imaging, sub-acute follow-up MRI, and clinical tabular data. The contributions of this work are two-fold: first, we introduce a standardized benchmarking of final stroke infarct segmentation algorithms through the ISLES’24 challenge; second, we provide insights into infarct segmentation using multimodal imaging and clinical data strategies by identifying outperforming methods on a finely curated dataset. The outputs of this challenge are anticipated to enhance clinical decision-making and improve patient outcome predictions. All ISLES’24 materials, including data, performance evaluation scripts, and leading algorithmic strategies, are available to the research community following https://isles-24.grand-challenge.org/.
keywords:
Ischemic stroke , Lesion segmentation , Final infarct , Multimodal data
journal: Elsevier
1 Introduction
Accurate segmentation of ischemic stroke lesions is crucial for guiding treatment decisions during acute stages (e.g., determining eligibility for thrombectomy), evaluating outcomes, clinical follow-up, and optimizing therapeutic strategies in sub-acute and chronic stages. The Ischemic Stroke Lesion Segmentation (ISLES) Challenge (https://www.isles-challenge.org/) significantly advances stroke image analysis by bringing together experts in neurointervention, radiology, and computer science from top institutions. Biomedical challenges like ISLES are now considered the de facto gold standard for algorithm comparison by the research community and have been organized multiple times, covering diverse pathologies and imaging modalities, thereby further cementing their role in advancing medical imaging technologies [1, 2, 3, 4, 5, 6].
The ISLES Challenge has been a recurring feature at MICCAI. The inaugural ISLES’15 focused on segmenting sub-acute ischemic stroke lesions from post-interventional MRI and acute perfusion lesions from pre-interventional MRI [7]. Subsequent challenges, ISLES’16 and ISLES’17, emphasized stroke outcome prediction by requiring the segmentation of follow-up stroke lesions from acute multimodal MR imaging and estimating patient disability scores [8]. ISLES’18 addressed acute stroke segmentation indirectly and cross-modally by predicting core tissue delineated in concomitant MRI from acute perfusion CT series [9]. The latest ISLES’22 expanded to segment acute, sub-acute, and chronic ischemic strokes in over 2000 MRI scans [10, 11, 12]. Thus, the ISLES Challenge has garnered considerable attention over the years.
Building on this rich history, ISLES’24 aims to segment the final stroke infarct using pre-interventional acute stroke data. Unlike standard clinical imaging techniques for core estimation, participants have access to the full CT trilogy (non-contrast CT (NCCT), CT angiography (CTA), and perfusion CT (CTP)); follow-up imaging data (DWI and ADC); and clinical tabular data (demographics, clinical history, laboratory results, and stroke outcomes). This creates a unique 360-degree challenge setting where all feasibly accessible clinical data are available for participants. Additionally, ISLES’24 enables the combination and learning from diverse CT techniques with and without contrast, evaluation of brain tissue changes before and after recanalization treatment through follow-up imaging, and inclusion of variables known to impact ischemic tissue status. This comprehensive approach allows participants to refine segmentation models through multimodal data learning strategies. In brief, the main contributions of this work are three-fold:
  1. 1.
    The standardized benchmarking of final, post-treatment stroke infarct segmentation algorithms through the ISLES’24 challenge (https://isles-24.grand-challenge.org/).
  2. 2.
    Performance evaluation of multimodal imaging and clinical data strategies over a finely curated stroke dataset, including clinical data, acute (NCCT, CTA, CTP), and follow-up (DWI, ADC) stroke imaging.
  3. 3.
    The open access provision of ISLES’24 materials for the research community, including data, performance evaluation scripts, and leading algorithmic strategies.
2 Related work
2.1 Final infarct prediction using perfusion CT
Perfusion CT is the reference imaging modality for predicting stroke lesion volumes in clinical settings. The workflow involves estimating perfusion maps using established deconvolution approaches and then deriving penumbra and core volumes through thresholding [13]. Although the preferred perfusion map for initial infart (core) estimation varies among software vendors, cerebral blood flow (CBF) and cerebral blood volume maps have been widely adopted. In the last years, various studies propose deep-learning methods for estimating stroke lesions. For instance, de la Rosa et al. [14] propose a deep-learning enhanced SVD-deconvolution method to estimate core tissue. Robben et al. [15] bypass traditional deconvolution by predicting final infarct volumes using native CT Perfusion images and clinical metadata. Their approach combines diverse inputs—CTP, time and treatment parameters, and CTP-derived vascular functions—into a DeepMedic-based model [16], allowing for the simulation of diverse clinical scenarios. Validation of their model over clinical data shows improved estimation of final infarct compared to traditional deconvolution software. Building on this foundation, Amador et al. [17, 18] introduce a deep spatio-temporal CNN to predict final infarct from raw perfusion data. Their model combines a U-Net-like architecture with temporal convolutional networks to efficiently process 4D data. de Vries et al. present PerfU-Net [19], a spatio-temporal model for estimating core tissue from CTP imaging data. PerfU-Net employs a U-Net-based architecture enhanced with an attention module, allowing it to capture dynamic perfusion patterns. Vision transformers have been introduced for infarct segmentation by de Vries et al. [20] and have been later combined with CNN models through hybrid methods in [21]. Gutierrez et al. [22] propose the prediction of follow-up CT images from CTP data using spatio-temporal models, from which the final infarct can be derived from the generated pseudo-CT images. More recently, by leveraging physics-informed neural networks, de Vries et al. [23, 24] show promising results while using a hybrid data-driven image modeling approach.
2.2 Beyond ASPECTS: Final infarct segmentation using non-contrast CT
Machine and deep learning strategies for infarct segmentation have increasingly been applied to native NCCT due to its broader availability compared to advanced imaging techniques like CTA and CTP. While NCCT reveals infarcted brain tissue as hypodense areas, these changes typically become visible later than those observed with perfusion imaging. Early ischemic changes on NCCT are challenging due to low signal-to-noise ratio, subtle signals, artifacts, and confounding factors [25]. In this context, Srivatsan et al. [26] proposed an image-processing technique for segmenting early ischemic changes by creating a relative NCCT map that highlights subtle hypoattenuation differences between contralateral hemispheres. Qiu et al. [27] trained a standard U-Net model to segment ischemia. Their model, validated over 100 independent scans, demonstrated high volumetric agreement with DWI-derived annotations. Similarly, El-Hariri et al. [28] assessed the performance of the self-adapting nnU-Net model, showcasing high volumetric agreement with ground truth and intraclass correlations over 80%. In 2021, Pan et al. [29] tested a ResNet-based CNN combined with a maximum a posteriori probability model. Later on, Kuang et al. introduced EIS-Net, a multi-task learning model designed to simultaneously segment early infarcts and score the Alberta Stroke Program Early CT Score (ASPECTS) [30]. EIS-Net employs a 3D triplet convolutional neural network that incorporates original and mirrored NCCT images along with an atlas to improve feature extraction. This model excelled in both segmentation and ASPECTS scoring tasks, showing high volumetric agreement with ground-truth annotations and accurate ASPECTS ratings. Wang et al. [31], instead, showed marginal gains of a SwinUNETR model compared to a traditional U-Net method. More recently, Kuang et al. [32] leveraged circular feature interaction and bilateral difference learning in a hybrid CNN-Transformer model to segment ischemia on NCCT.
2.3 CTA-based final infarct prediction
CTA is a vital imaging modality in stroke management, primarily used to assess cerebral vasculature and identify large vessel occlusions (LVOs). While CTA is traditionally employed for diagnostic purposes, several studies have revealed the potential of CTA-derived features in infarct prediction [33, 34]. Unlike perfusion CT, which provides direct information about blood flow and tissue viability, CTA-based approaches rely on vascular biomarkers such as collateral flow [35, 36, 37] and ASPECTS [33, 38, 36] to estimate the final infarct size.
Building on these findings, Sheth et al.[39] developed an automated deep learning model called DeepSymNet to identify large vessel occlusions (LVO) and the infarct core from CTA source images. Their method demonstrated an ability to determine the infarct core with an AUC of approximately 0.88. Recently, Hokkinen et al.[40] reported that a CTA-based CNN method showed a moderate correlation with final infarct volumes in the late time window for patients with large vessel occlusion who were successfully treated with endovascular therapy. Giancardo et al. [41] further advanced the field by introducing a deep learning method that incorporates a novel weighted gradient-based approach for stroke ischemic core segmentation using image-level labeling. This approach outperforms nnUNet models trained on voxel-level data. More recently, Palsson et al. [42] introduced a U-Net-based convolutional neural network for lesion prediction, augmented with a spatial and channel-wise squeeze-and-excitation block. Their study demonstrates the feasibility of predicting 24-hour follow-up infarcts using acute CTA imaging, achieving results comparable to those of CTP-based algorithms.
2.4 Multimodal final infarct prediction models
While single-modality approaches using CTP, NCCT, or CTA have shown promise in predicting final infarct volumes, there are growing research efforts on combining multiple imaging modalities for enhanced accuracy and reliability in final infarct prediction. Multimodal approaches could leverage the complementary strengths of different imaging techniques, potentially enabling a more comprehensive assessment of ischemic stroke. In this context, Wang et al. [43] developed a hybrid multiscale CNN model that integrates NCCT, CTA, and CTA+ (acquired with an 8-second delay after CTA) to identify the infarct core and deficit. Their quantitative analysis demonstrated a high level of agreement with CTP-based definitions, highlighting the feasibility of employing commonly accessible imaging modalities (NCCT and CTA) for ischemic stroke assessment using deep learning. Other studies also explored the benefits of combining imaging data with tabular clinical information for enhanced stroke prediction. For instance, Robben et al.[15] leverage models including images, vascular functions, and treatment information. Building on this approach, Amador et al.[44] developed a multimodal spatio-temporal model that employs cross-attention mechanisms to simultaneously fuse information from 4D CTP and clinical metadata, resulting in improved accuracy in stroke infarct prediction.
3 Methods
Dataset
The dataset used in ISLES’24 has been specially prepared for the challenge. It includes multi-scanner and multi-center data derived from large vessel occlusion ischemic stroke cohorts. The dataset includes acute and sub-acute stroke imaging and clinical (tabular) data. Acute imaging data has been acquired at patient admission and includes the diagnostic CT trilogy: NCCT, CTA, and CTP, as well as CTP-derived perfusion maps (namely CBF, cerebral blood volume, mean transit time, and time-to-maximum a.k.a Tmax). Follow-up imaging data was acquired 2 to 9 days later and included DWI and ADC. Moreover, the clinical data contains demographics (age, sex, etc.), patient history, admission scores (e.g., admission NIHSS), and clinical outcomes (e.g., 3-month mRS). In Figure XXX, an example data case is shown.
The dataset is released in raw and preprocessed formats, thus allowing participants to devise algorithms with diverse degrees of freedom. For anonymization purposes, all scans are defaced based on brain and face masks obtained with TotalSegmentator [45], and the clinical tabular data de-identified by adding $\pm$ 5% random noise to the laboratory variables. Preprocessing of the CTP series has been performed using the FDA-approved clinical software icobrain cva [46, 47]. The 4D CTP series are motion-corrected through image co-registration and temporally resampled at 1 frame/second. Afterwards, perfusion maps are generated using a traditional tracer-kinetics deconvolution algorithm. The CTA scans are released after cropping the images around the head, thus removing the thoracic region when available. Preprocessing of the images has been performed by linearly interpolating and registering all the imaging series to the NCCT scans. Except for the MRI scans, where affine transformations are used, all remaining images are registered following rigid transformations. Registration is performed using the Elastix [48] and NiftyReg [49] toolboxes. Furthermore, MRI scans are skull-stripped using HD-BET [50]. Lesion masks are derived from DWI images using the ISLES’22 ensemble algorithm [12]. Quality control and correction of the lesion masks are performed when needed by medical students (TAB, HPM) supervised by two neuroradiologists (JSK, BW) with more than 10 years of experience.
The dataset (N = 250) is split into train (N = 150) and test subsets (N = 100). The train (test) set contains N = 100 (N = 50) scans from the University Hospital of Munich and N = 50 (N = 50) scans from the University Hospital of Zurich. All the train subset is publicly available, while the test subset is hidden from the public, precluding participants from model overfitting strategies. The ISLES’24 data is structured and maintained following the Brain Imaging Data Structure (BIDS) guidelines [51].
Figure 1: Overview of the ISLES’24 task and data modalities. Participants can exploit 1D/2D/3D/4D final infarct learning strategies. NCCT: non-contrast CT; CT: CT angiography; CTP: perfusion CT; CBV: cerebral blood volume; CBF: cerebral blood flow; MTT: mean transit time; Tmax: time-to-maximum; DWI: diffusion weighted imaging.
3.1 Task
The ISLES’24 challenge task is the segmentation of final, post-treatment infarct brain tissue from pre-treatment imaging and clinical data. Participants can choose which image modalities to use and if using them raw or already preprocessed. Multimodal data approaches can be used, as the data involves 1D, 3D, and 4D modalities. An overview of the challenge task and the involved data modalities is depicted in Figure 1.
Algorithms’ evaluation
Metrics
[…]
Ranking
The final challenge ranking is determined by evaluating each method’s performance on the unseen test set. The test phase of the challenge is blind and single-shot; participants do not have access to the input images, their output results, or the ground truth images. This setup entirely precludes any manual intervention, visual quality assessments, or case-specific overfitting strategies by the teams.
The ranking is performed as in previous ISLES editions [7, 8, 9, 12]. The ranking scheme is derived from the average ranking position obtained from all the individual rankings by case and by metric. First, for each scan and metric, the participants are ranked. Then, for each participant, the average rank across all the metrics for each case is computed, leading to a ’per-case average rank’ for each team. The final ranking is obtained by averaging these ’per-case average ranks’ across the entire dataset.
Data and Code availability
The ISLES’24 dataset is accessible at at https://isles-24.grand-challenge.org/. To facilitate team participation and metric quantification, two Git repositories have been pre-released. The first repository, https://github.com/ezequieldlrosa/isles24, provides an introduction to the challenge and includes a Python notebook that demonstrates the computation of performance metrics. The second repository (https://github.com/ezequieldlrosa/isles24-docker-template) offers a template Docker algorithm, designed to streamline the creation of algorithmic Dockers. The F1-score and absolute lesion count difference metrics are computed using the panoptica Python library [53]. Connected components are identified using the cc3d [54] Python package. The final challenge ranking is determined using the rankThenAggregate function from the challengeR [55] package implemented in R [56].
4 Results
This section presents the ISLES’24 algorithmic rankings and performance metrics obtained during the hidden test phase. For the best-performing methods, we also conduct further analysis to gain deeper insights into their performance, encompassing both quantitative and qualitative evaluations.
5 Discussion
Predicting final infarct stroke lesions from pre-interventional acute data is a complex task due to multiple factors. Acute CT imaging, while widely used, is less sensitive compared to the gold-standard DWI and often suffers from artifacts (e.g. streak) and low signal-to-noise ratio. These technical limitations make it particularly challenging to identify small stroke lesions, such as embolic shower patterns, which are crucial for understanding the full extent of the stroke (Boers et al., 2013). Furthermore, the dynamic nature of stroke progression adds additional complexities. Infarct growth can occur unpredictably, and small DWI lesions may even be generated during the recanalization procedure itself (Albers et al., 2006). Edema can compress brain structures, altering anatomy in ways that may not be visible in acute images (Kucinski et al., 2002). Additionally, technical challenges in data processing and handling, such as image interpolation and registration due to different image spaces and resolutions between acute CT and follow-up MRI, further complicate accurate segmentation (Crum et al., 2004). This multifaceted problem underscores the need for advanced techniques and careful interpretation in stroke lesion segmentation.
Acknowledgement
SW is supported by the Swiss National Science Foundation (Grant No. 310030_200703), the UZH Clinical Research Priority Program (CRPP) Stroke, and the Swiss Heart Foundation. EdlR and BM are supported by the Helmut Horten Foundation. EOR is supported by the TUM KKF Clinician Scientist Program